the length of a crayon box base is 3 more than 4 times its width.IF the perimeter of the crayon box is 21 centimeters,find the lenght and width.
Answer by checkley77(12569)

L=3+4W
P=2L+2W
21=2(3+4W)+2W
21=6+8W+2W
21=6+10W
10W=21-6
10W=15
W=15/10
W=1.5 C. ANS.
L=3+4*1.5
L=3+6
L=9 C. ANS.
PROOF:
21=2*9+2*1.5
21=18+3
21=21